MARS DESERT RESEARCH STATION

Astronomy Report – March 1

ASTRONOMY REPORT
SOL 17

 

NAME:  Mouadh Bouayad            CREW: 175

DATE: 03/01/2017

DATE OF OSERVATION: 02/28/2017

SKY CONDITIONS: clear
WIND CONDITIONS: no wind

OBSERVATION START TIME: 8:00 pm
OBSERVATION END TIME: 12:00 am
SUMMARY: Yesterday afternoon, I took about an hour to become familiar with the CCD camera, and the software on the astronomy computer. I prepared everything to be able to use it right after (when night falls). I also figured I could use the software from the hab, as I found the blue Ethernet cable in the hab. It finally took me more time than I expected to do the focus, and I therefore missed Mars and Venus. I however could take some more pictures of Orion Nebulae and Jupiter.

 

Nevertheless, I still don’t know how to zoom in or out, especially in order to get more details of Jupiter. Because no matter how hard I tried to dim it down, it still seemed too bright to get some details. Now that I know more about how everything works, I think I’ll be able to take a lot more pictures tonight.

 
OBJECTS VIEWED: Orion Nebulae, Betelgeuse (no photos), Jupiter
PROBLEMS ENCOUNTERED: None

Science Report – February 28th

Science Report

Experiment : Optinvent AR Glasses
Person filling in the report: Louis Maller

I have tinkered some more with Tasker and have therefore managed to implement an interesting function on the glasses:
When you tilt the head to the left, it starts recording what you are saying, and when you tilt your head to the right it stops recording. This functionality was tested successfully during today’s EVA, with the help of Louis Mangin. We have a recording of him giving some information (level of the water tank or propane for example). The voice is clear and understandable even inside the helmet.

The tilt of the head needed is quite strong, the glasses need to be brought to a vertical position, so you have to tilt the body somewhat in order for this to work. I’m currently working to solve this issue.

During the EVA, the glasses and the computer were both connected to the same LAN network created by my phone.  The phone was on the second floor of the hab. I could see on my screen exactly what the person wearing the glasses could see, though with a slight lag. The connection was lost as they went to check the propane tank, when they disappeared behind the small hill. The disappointing thing was that even when they came back close to the hab, I was unable to get the connection with the glasses. I am also trying to solve this issue. It seems to require relaunching the AirDroid App that I am using, which I can also do by using position of the head (for example head tilted upwards works).

Experiment: Seismometer
Person filling in the report: Mouadh Bouyad

I have read the data we collected yesterday, and I think that we have coherent data, especially when it comes to the temperature. I have however noticed something weird : in most of daily measurements, I
obtain irregularities that seem very unnatural to me, almost at the same time (4pm french time).
I wonder if they are not data due to the fact that we sometimes change the key ; however, I can see the same kind of anomaly on other days, when measurements have been made continuously (without changing the key).

Science Report – February 27th

Science report

Experiment: Optinvent AR glasses
Person filling in the report: Louis Maller
The glasses were taken out today once again. They properly functioned during the entire engineering check (part of the team that stayed near the hab, not the one that went all the way to the diesel tank). The signal was lost only when I drove off on the ATV. That’s a good success.
Optinvent gave me access to some software that would allow me to control my glasses from my phone, but it still requires some troubleshooting because of Bluetooth issues.
Also I have been tinkering with Tasker to try to turn on apps with movements from the accelerometer, but I’m not an expert and need to improve what I have started doing.

Experiment: Seismometer
Person filling in the report: Mouadh Bouayad
We went to get back the data today; there was a lot of wind. We even had to put the sensor back to  its place again but it had barely moved. Otherwise everything was ok today.

Experiment: Aquapad:
Person filling report: Arthur
Yesterday I put a few drops of filtered water and boiled water in two
aquapad petri dishes and let them at 35°C in the Science Dome oven.
This afternoon (24 hours after), I took them out of the oven and used
the iPad application “Everywear” developed by the CNES, to count the
red points in the Aquapads (the points are caused by the growth of the
bacteria present in the water). Unfortunately, the glass of the petri
dishes was covered with fog, but it was possible to conclude that the
bacterial pollution of our water is stable: I did not see any major
variation in the quantity of red points, compared with mo observations
of the two last weeks.

Experiment: Sextant:
Person filling report: Arthur
I used the sextant three times during today’s EVA, but lost the paper
map because of the strong wind. As a result, in the afternoon I
figured out a way to determine my position on the PDF map by drawing
the lines and angles on a graphics software.
I join to this report the part of the map showing the exact position
of the seismometer, South of the Hab, with additional information
about the sextant angles.

MDRS seismo position

Astronomy Report – February 27th

ASTRONOMY REPORT

SOL 15

NAME:  Mouadh Bouayad            CREW: 175

DATE: 02/27/2017

DATE OF OBSERVATION: 02/26/2017
SKY CONDITIONS: mostly clear, few clouds here and there.
WIND CONDITIONS: no wind
OBSERVATION START TIME: 8:00 pm
OBSERVATION END TIME: 12:00 am
SUMMARY: Arthur and I went to the observatory to observe Andromeda Galaxy, and took few pictures as well. We then could take some more pictures of Orion Nebulae, and then Jupiter. The images are clearer! As you can see on the photos. We couldn’t however have a better focus on Jupiter, neither could we zoom in.
OBJECTS VIEWED: Andromeda Galaxy, Orion Nebulae, Jupiter.
PROBLEMS ENCOUNTERED: None

 

 

Astronomy Report – February 26th

ASTRONOMY REPORT

SOL 13

 

NAME:  Mouadh Bouayad            CREW: 175

DATE: 02/25/2017

SKY CONDITIONS: mostly clear, few clouds here and there.
WIND CONDITIONS: a little bit of wind
OBSERVATION START TIME: 8:30 pm
OBSERVATION END TIME: 11:45 pm
SUMMARY: We finally had favorable weather conditions yesterday. Arthur and I first went there in order to observe Mars and Venus. However, by the time we got there, the two planets had already disappeared from the sky. Therefore, we observed Betelgeuse, Sirius, and Orion Nebulae. We took pictures of each (see below). We finally observed and took a photograph of Jupiter, but it was too bright, and we it was too cold to try and get a better photograph. So we went back to Hab at 11:45 pm.

I know we should have taken about 50 photos of each objects, in order to have as most photographs as possible, in order to make an image treatment and the better of them all.
OBJECTS VIEWED: Betelgeuse, Sirius, Orion Nebulae, Jupiter
PROBLEMS ENCOUNTERED: None, except for the cold (about 24°F).

Sirius

 

Jupiter

 

Bételgeuse

 

Nebuleuse Orion

Science Report – February 24th

Science Report

Experiment : Optinvent AR Glasses
Person filling in the report: Louis Maller

Finally found an application that allows to connect the glasses to a computer using a local ad-hoc network (LAN). The HabCom was able to view on his computer what I was seeing. Unfortunately, the diesel tank was already out of range for the LAN. The phone creating the LAN was in my pocket, maybe placing it inside the GreenHab for example would have allowed for better coverage. Tomorrow we will do some range tests to evaluate this possibility during the EVA. Further work will be based on trying to find a solution to control the glasses (operating on android) from my phone (also on android), through a LAN. Suggestions welcome!

Experiment: Sextant
Person filling in the report: Arthur Lillo
The Sextant has proved that it is usable during EVAs even with gloves. Thanks to the new compass that was built, it is much more easy to find our position on the map. The development of a dedicated app is on standby due to lack of appropriate software.

Experiment: Balloon
Person filling in the report: Simon Bouriat
There was a deployment test this morning, which was unsuccessful due to strong gusts of wind that took us a bit by surprise. We aborted the deployment, and as we tried to put the balloon back in place, it ripped a little in some places. Xavier and I spend some time this afternoon fixing it. We are hoping that the wind will lower enough before the end of our mission in order for us to deploy it successfully, ideally for 24 hours.

Experiment: Aquapad
Person filling in the report: Arthur Lillo
Aquapads were prepared today in order to test the evolution of the quality of the filtered water in the Hab.

Science Report – February 19th

Science Report 19Feb2017

Experiment: Navigating on Mars
Person filling in the report: Arthur Lillo

I took the sextant during the EVA: it was easy to use and to read
despite the gloves and the fog in the helmet. However, the temperature
gradient decalibrated the mirror before long, and the calibration
screws are too small to be turned with the gloves. Therefore, I could
only take one measurement: I wrote down the angle between the Hab and
Marble Ritual, and the angle between Phobos Peak and Marble Ritual.
Back inside the Hab, I drew on the map the two arcs of a circle
corresponding to the two angles, and I obtained a point that seems to
be a few meters away from my position during the measurement. Thus,
the experiment was a success, now I need to code an app that does the
work of positioning.

Experiment: EVA Emergency Procedures
Person filling in the report : Simon Bouriat

Today was the first line of our abstract about emergency procedures on EVA. The idea is to build emergency procedures based on a study of the ones of mountaineering expeditions. To do that, we want to simulate, during our EVA, different emergency situations. Nowadays, such situations are not currently studied.
One member will make a planning to prepare two or three simulations during the next two weeks. No other member, except for the one that will simulate the incident, will know that an issue will occur. But before, we want to study and create emergency procedures that will be taught to the all crew. The simulations will be a way to apply these procedures.
This study will probably continue after the sim.

Since we had no PRIF for this experiment, we would like to know if we can carry it out.

Science Report – February 18th

Science report

Experiment: Seismometer

Person filling in the report: Mouâdh Bouayad

We went to check the seismometer during today’s EVA; everything was just perfectly fine. Even the level of the seismometer didn’t move at all. We changed the USB key to read the collected data. I could analyze them by the time we got back to the Hab, and no problem. Unfortunately, I have a software pretty hard to use to analyze data. Actually, the digitizer cuts the data every hour. Thus, in order to view the signal on a long period of one axis, we need to merge the curves, which I don’t yet know how to do with the software (named PQL) that I have installed.

Other experiments ongoing (Aquapad analysis continued, work on the Balloon equipment, some tests with the AR glasses).

Science Report – February 17th

Crew 175 Science Report 17Feb2017

Science Report SOL 5

Experiment: Seismometer
Person filling in the report: Mouadh Bouayad

We could deploy the seismometer this morning, during the EVA. It was quite hard to do, with the gloves, especially to put the seismometer correctly. It was nevertheless a success. We decided to put the Stanéo digitizer in a box in order to protect it, and we covered it with a piece of plastic. We have then tried to immobilize everything by putting stones here and there.
The weather on the rest of the sol was worrying: there was a lot of wind (impossible thing on Mars because of the low pressure on its surface). I fear that the structure might crumble. Unfortunately, we’ll have to wait until tomorrow to check this out, during the next EVA.

Experiment: Balloon
Person filling in the report: Simon Bouriat

We were first supposed to use the balloon during the EVA this morning. Unfortunately, the weather was not on our side. We still don’t have enough results to make any conclusions. We hope to use it again next week.
The crew engineer and I started working on our new project about emergency procedures. There are three main points in that study. First, we will study the emergency procedures observed during scuba diving. One of the big step in this study is the comparison between this sport’s dangers and EVA’s ones. Secondly, we will try to set up different emergency procedures depending on the issues and brief the rest of the crew about it. Finally, we will simulate unexpected health or material issues during EVAs and try to apply these procedures.

Simon Bouriat – HSO Crew 175

Science Report – February 16th

Science report

Experiment: Balloon
Person filling in the report: Simon Bouriat
The balloon went on the EVA yesterday. It worked really well even if the weather was a bit windy. Unfortunately, the platform broke when we stored it. Today, during the afternoon, I fixed it and I recoded the Arduino. The idea is to record pressure and temperature in the hab during the whole day tomorrow and to compare the results with the ones we got from the balloon flight (here after). It will give us the average characteristics on the ground.  We also want to calculate the height of the balloon thanks to our results. Tomorrow, during the EVA, we will try to tether the balloon with two ropes to avoid any rotation of the new platform.
I also tried to figure out how to build a radio relay. The main way in my opinion is to use one of the broken walkie-talkies.



Experiment: Seismometer
Person filling in the report: Mouâdh Bouayad
Today the battery was wrapped in an emergency blanket in order for it to be protected, especially from water. The assembly of the science equipment (ISAE-SUPAERO Seismometer and Staneo acquisition system) was assembled with the gloves in order to train for tomorrow’s EVA.

Experiment: Aquapad
Person filling in the report Arthur: Lillo
The Aquapad (CNES) experiment has started today, with water samples having been prepared for testing, it has been put in the science oven, the first results will be available tomorrow!

Experiment: EMUI
Person filling in the report: Louis Maller
During the EVA I wore the glasses to test different issues. Comfort: they are lighter and a lot less trying to wear than the previous version, this is a true improvement. The right branch of the glasses still hurt a bit, at the end of the EVA I was glad to take them off, maybe because of the cable attached to the external battery pulling down?
The glasses were attached to an external battery and I went through the entire EVA without any problems.
I was able to see the screen only if I managed to make some shade on the glasses, or if I was in a dark place, like in the shadows. The time widget was very visible.
Screencasting is still an issue for now as the available software either require internet to work, or require a cable…